# Controllable diatomic molecular quantum thermodynamic machines

### v1

@@ Abstract
We present quantum heat machines using a diatomic molecule modelled by a $q$-deformed potential as a working medium. We analyze the effect of the deformation parameter and other potential parameters on the work output and efficiency of the quantum Otto and quantum Carnot heat cycles. Furthermore, we derive the analytical expressions of work and efficiency as a function of these parameters. Interestingly, our system reveals different operational phases for the range of parameter regimes considered, such as a quantum heat engine. In addition, the efficiency of the quantum Otto heat engine is seen to be tunable by the deformation parameter. Our findings provide useful insight for understanding the impact of anharmonicity on the design of quantum thermal machines.
@@ I Introduction
Quantum heat engines produce work using quantum matter as a working substance and exhibit exotic properties. They have been shown to perform better compared to their classical counterpart [3]. Several quantum mechanical models have been proposed as a working medium in recent times, for example, the harmonic oscillator potential [4], infinite square well potential [4], spin systems [5, 6, 7, 8], stanene [9], strained-graphene [10], Dirac particles [11, 12], two-level system [13], multi-level system [14], a continuum working medium [15], photon gas [16], diatomic molecules [17], Posch-Teller potential [18, 19], among others. In addition, the past decade has witnessed several experimental implementations of quantum thermal machines [20, 3, 21, 22, 23, 24, 25].
@@
Moreover, quantum systems that exhibit anharmonicity, such as molecular systems [26, 27], superconducting qubits [28], are of great interest in the development of quantum technologies. Anharmonic effects have been shown to be useful in several settings [29], but most quantum systems are often treated with approximate harmonic theoretical descriptions, for instance in the case of a single molecule [30, 31, 32, 33, 34, 35, 36, 37]. The harmonic approximation is valid when atomic displacements are rather limited [38], and recent experiments are beginning to probe the limits of this model in a plethora of processes and contexts [39, 40, 41]. However, most quantum thermal machines assume that their working mediums are often harmonic in nature [42, 20, 3], but recent experiments show the possibility of trapping molecules in anharmonic traps [29]. Therefore, it becomes crucial to develop quantum thermal machines with genuine anharmonicity to see how they impact the performance of thermal machines. In this regard, a Carnot-like engine with a diatomic molecule has recently been proposed by Oladimeji et al. [17]. Here, we extend this to carry out a full quantum thermodynamic description of a quantum thermal machine whose working medium is a tunable Morse potential, which gives us the “degree of freedom” to tune the behaviour of the model. The Morse potential is a well-known diatomic molecular potential that captures the effect of anharmonicity and real bond-breaking [26]. We investigate the anharmonic effects on the performance of a quantum thermal machine in an analytically feasible model using a generalized ($q$-deformed) Morse potential as the working substance. The choice of this model is inspired by the fact the presence of the $q$-parameter gives us tunable control in our setup. This $q$-deformed physical system is typically described by q-deformed potentials which is reduced to their normal equivalent when $q\rightarrow 1$. The q-deformed hyperbolic potentials first appeared in the seminal works of Arai [43, 44]. This class of model potentials has since found applications in several areas of research, notably the study of the phonon spectrum in ${}^{4}\text{He}$ [45], multi-atomic molecules [46], oscillatory-rotational spectra of diatomic [47], the description of the electronic conductance in disordered metals and doped semiconductors [48] amongst many other[49, 50, 51]. A recent study showed that statistical complexity on the information-theoretic measure of diatomic molecules is strongly influenced by the potential deformation parameter $q$ [52]. Therefore, it is the goal of this study to extend the application of such a physical system to quantum thermodynamics.
@@ II The diatomic molecule
Figure 1: Plot of the q-deformed Morse potential for $q\!\in\!\{0.4r_{e},~{}0.5r_{e},~{}r_{e}\}$, where $q\rightarrow 1$ corresponds to the standard Morse potential. Parameters used: $\mathcal{D}_{\text{e}}\!=\!10r_{e}$, and $\alpha\!=\!2r_{e}$.
@@
It is interesting to note that in contrast to the standard Morse potential, $q$-deformed Morse potential cannot be approximated by the quantum harmonic oscillator except in the limit of $q\rightarrow 1$, $\mathcal{V}_{q}(x)\!\approx\!-D_{e}+\xi^{2}D_{e}x^{2}+O\left(x^{3}\right)$ in the neighbourhood of its minimum. Hence, we can see that for small values of $x$ the Morse potential, Eq. (1), can be approximated by a harmonic oscillator with spring constant $k=2\xi^{2}D_{e}$, and $V_{q}(x)\!\approx\!-D_{e}+k/2x^{2}$, this was illustrated by Leonard and Deffner [56]. It is observed that the Morse potential significantly deviates from the harmonic approximation as the distance from $x\!=\!0$ increases. This allows us to model bond weakening and dissociation with diverging nuclear separation. Moreover, the presence of the deformation reveals other features such as symmetry distortion [52]. This is particularly evident in the regime where the $q$ parameter is very small. The consequence of these properties introduces interesting features into the quantum heat engine performance, as we will discuss shortly in Sec. III.
@@
where $\beta\!=\!1/(k_{\mathrm{B}}T)$ is the inverse temperature $T$ with the Boltzmann constant $k_{\mathrm{B}}$. In the energy representation, it is straightforward to calculate the partition function using the $q$-deformed Morse potential eigen energies (3), we obtain
@@ III.1 Quantum Carnot Cycle
[…]
@@
[…]
@@
Figure 2: A pictorial representation of a Carnot thermal machine using a $q$-deformed anharmonic oscillator as a working substance. The engine cycle consists of two adiabatic strokes ($B\to C$ and $D\to A$) and two isothermal strokes, $A\to B$ at $T_{\text{h}}$ and $C\to D$ at $T_{\text{c}}$, with $T_{\text{h}}>T_{\text{c}}$.
Figure 3: Quantum Carnot engine: (a) The heat exchanges of the working substance with the hot reservoir, $\mathcal{Q}_{\text{h}}$ vs $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$ (b) The heat exchanges of the working substance with the cold reservoir, $\mathcal{Q}_{\text{c}}$ vs $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. (c) The Work output $\mathcal{W}$ vs $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. Parameters used: $\alpha_{h}=2.236r_{e}$, $\alpha_{c}=r_{e}$, $T_{c}\!=\!2r_{e}$ and $T_{h}=10r_{e}$
@@
In Fig. (3), we present the numerical results of the heat exchanges and the total work extracted against the control/deformation parameter ($q/r_{e}$) and the dissociation energy ($D_{e}/r_{e}$). Fig. 3 (a) shows that the input heat is positive, that is $\mathcal{Q}_{\text{h}}>0$ while Fig. 3(b) demonstrate that the output heat is always negative $\mathcal{Q}_{\text{c}}<0$ within the parameters range. In Fig. 3(c), we observe an increase in work output as both $q/r_{e}$ and dissociation energy $D_{e}/r_{e}$ increase. Both parameters enhance the work output but have a stronger effect when they are at their maximum. This demonstrates that the width of the well and the hot bath temperature play a crucial role in enhancing the work output. Thus, considering that the total work output is positive, we conclude that the system operates as a quantum heat engine in this regime. The efficiency of the quantum Carnot $q$-deformed Morse oscillator heat engine is 0.8 at temperatures that satisfy the reversibility condition.
@@ III.2 Quantum Otto Cycle
[…]
@@
[…]
@@ III.2.1 Changing width of the potential well
where the dimensionless parameters $\Lambda_{2}^{\text{hc*}}2\gamma_{1}^{\text{c}},\gamma_{1}^{\text{h}},\text{and}\Lambda_{4}^{ch}$ are defined explicitly in Eq. (35), and (43) of Appendix (B). Using Eq. (17), we obtain the work output as follows:
$\mathcal{W}^{\text{QO}}=\frac{1}{2\sqrt{\pi}}\left(\frac{\Lambda_{10}+\Lambda_{11}\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}}{\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}\text{{\hbox{\beta}}c}\beta_{\text{h}}\xi_{\text{h}}^{2}\text{erfc}\left(\gamma_{1}^{\text{h}}\right)}-\frac{p\mathcal{R}_{0}}{\mathcal{R}_{1}}\right)$ (21)
@@
Figure 5: Quantum Otto engine with changing width of the potential well: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$, (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ (c) Work output, $\mathcal{W}$, and (d) Efficiency, $\eta$ as a function of $\mathcal{D}_{e}/r_{e}$ and $q/r_{e}$. Parameters used: $\alpha_{h}\!=\!2.236\,r_{e}$, $\alpha_{c}\!=\!r_{e}$, $T_{c}\!=\!2\,r_{e}$ and $T_{h}\!=\!10\,r_{e}$.
@@
Figure 6: Quantum Otto cycle with the changing deformation parameter: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$ (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ (c) Total work output, $\mathcal{W}$ (d) Efficiency $\eta$ as function of $\mathcal{D}_{e}/r_{e}$ and $\alpha/r_{e}$. Parameters used: $q_{c}=0.8r_{e}$, $q_{h}=r_{e}$, $T_{h}=10r_{e}$, and $T_{c}=2r_{e}$.
@@ III.2.3 Changing the strength of the dissociation energy
Figure 7: Quantum Otto cycle for varying dissociation energy strength: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$ vs $\alpha/r_{e}$ and $q/r_{e}$. (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ vs $\alpha/r_{e}$ and $q/r_{e}$. (c) Work output of Otto engine $\mathcal{W}$ vs $\alpha/r_{e}$ and $q/r_{e}$. (d) Efficiency $\eta$ vs $\alpha/r_{e}$ and $q/r_{e}$. Parameters used:$\mathcal{D}_{e}^{\text{c}}=5r_{e}$, $\mathcal{D}_{e}^{\text{h}}=10r_{e}$, $T_{h}=10r_{e}$, and $T_{c}=2r_{e}$.
@@ IV Conclusions
We presented a theoretical proposal for a quantum thermal machine with the $q$-deformed Morse potential as a working medium. We discussed in detail the different regimes of operation for our heat engine. These modes of operation can be modulated by appropriate adjustment of the anharmonic potential parameters. Furthermore, we present analytical expressions for the work done and the efficiency of the heat engine. We find that these figures of merits depend on the variations of the potential parameters. It is also noted that the performance of both the engine and the refrigerator can be regulated or controlled by adjusting the deformation parameter and other system parameters. The potential platform for the realization of our model is a laser-cooled trapped ion as a microscopic heat machine, proposed by Gelbwaser-Klimovsky et al. [64]. However, in our context, the trap will be anharmonic, which will be more suitable for trapping a diatomic molecule. The experimental considerations for the realization of a tunable anharmonic trap have been presented by Home et al. [29]. Our findings may inspire several developments in the emerging field of quantum technologies such as the design of quantum thermal machines based on anharmonic models.

### v2

@@ Abstract
We present quantum heat machines using a diatomic molecule modelled by a $q$-deformed potential as a working medium. We analyze the effect of the deformation parameter and other potential parameters on the work output and efficiency of the quantum Otto and quantum Carnot heat cycles. Furthermore, we derive the analytical expressions of work and efficiency as a function of these parameters. Interestingly, our system operates as a quantum heat engine across the range of parameters considered. In addition, the efficiency of the quantum Otto heat engine is seen to be tunable by the deformation parameter. Our findings provide useful insight for understanding the impact of anharmonicity on the design of quantum thermal machines.
@@ I Introduction
Quantum heat engines produce work using quantum matter as a working substance and exhibit exotic properties. They have been shown to perform better compared to their classical counterpart [3]. In recent times, several quantum mechanical models have been proposed as a working medium, for example, the harmonic oscillator potential [4], infinite square well potential [4], spin systems [5, 6, 7, 8], stanene [9], strained-graphene [10], Dirac particles [11, 12], two-level system [13], multi-level system [14], a continuum working medium [15], photon gas [16], diatomic molecules [17], Posch-Teller potential [18, 19], among others. In addition, the past decade has witnessed several experimental implementations of quantum thermal machines [20, 3, 21, 22, 23, 24, 25].
@@
Moreover, quantum systems that exhibit anharmonicity, such as molecular systems [26, 27], superconducting qubits [28], are of great interest in the development of quantum technologies. Anharmonic effects have been shown to be useful in several settings [29], but most quantum systems are often treated with approximate harmonic theoretical descriptions, for instance in the case of a single molecule [30, 31, 32, 33, 34, 35, 36, 37]. The harmonic approximation is valid when atomic displacements are limited [38], and recent experiments are beginning to probe the limits of this model in a plethora of processes and contexts [39, 40, 41]. However, most quantum thermal machines assume that their working mediums are often harmonic in nature [42, 20, 3], but recent experiments show the possibility of trapping molecules in anharmonic traps [29]. Therefore, it becomes crucial to develop quantum thermal machines with genuine anharmonicity to see how they impact the performance of thermal machines. In this regard, a Carnot-like engine with a diatomic molecule has recently been proposed by Oladimeji et al. [17]. Here, we extend this to carry out a full quantum thermodynamic description of a quantum thermal machine whose working medium is a tunable Morse potential, which gives us the “degree of freedom” to tune the behaviour of the model. The Morse potential is a well-known diatomic molecular potential that captures the effect of anharmonicity and real bond-breaking [26]. We investigate the anharmonic effects on the performance of a quantum thermal machine in an analytically feasible model using a generalized ($q$-deformed) Morse potential as the working substance. The choice of this model is inspired by the fact the presence of the $q$-parameter gives us tunable control in our setup. This $q$-deformed physical system is typically described by $q$-deformed potentials which is reduced to their normal equivalent when $q\rightarrow 1$. The $q$-deformed hyperbolic potentials first appeared in the seminal works of Arai [43, 44]. This class of model potentials has since found applications in several areas of research, notably the study of the phonon spectrum in $^{4}\text{He}$ [45], multi-atomic molecules [46], oscillatory-rotational spectra of diatomic [47], the description of the electronic conductance in disordered metals and doped semiconductors [48] amongst many other[49, 50, 51]. A recent study showed that statistical complexity on the information-theoretic measure of diatomic molecules is strongly influenced by the potential deformation parameter $q$ [52]. Previous studies of quantum heat engines with Morse potentials have focused on the undeformed case [17]. Separately, q-deformed oscillators have been explored as effective models for anharmonic quantum systems. In contrast, the present work combines these two directions by introducing a q-deformed Morse potential as the working medium of a quantum heat engine. The deformation parameter $q$ provides an additional control knob that modifies the vibrational spectrum, enabling tunable a quantum heat engine.
@@ II The diatomic molecule
Figure 1: Plot of the q-deformed Morse potential for $q\!\in\!\{0.4,~0.5,~1\}$, where $q\rightarrow 1$ corresponds to the standard Morse potential. Parameters used: $\mathcal{D}_{\text{e}}\!=\!10$, and $\alpha\!=\!2$.
@@
For $q\neq 1$, the $q$-deformed Morse potential attains its minimum at $x=x_{0}=\ln(q)/\xi$ rather than at $x=0$. As a result, a Taylor expansion about $x=0$ contains a nonzero linear term and does not correspond to a harmonic approximation, whereas expanding about the true minimum $x_{0}$ yields a quadratic leading-order term and thus a well-defined harmonic approximation. Only in the limit $q\to 1$ does $x_{0}\to 0$, so that the expansion about $x=0$ coincides with the harmonic approximation. Hence, we can see that for small values of $x$ the Morse potential, Eq. (1), can be approximated by a harmonic oscillator with spring constant $k=2\xi^{2}D_{e}$, and $V_{q}(x)\!\approx\!-D_{e}+k/2x^{2}$, this was illustrated by Leonard and Deffner [56]. Although the Morse potential admits continuum states corresponding to dissociation, in this work we focus on its bound eigenstates, which already encode bond weakening through anharmonicity and the existence of a highest bound level. Moreover, the presence of the deformation reveals other features such as symmetry distortion [52]. This is particularly evident in the regime where the $q$ parameter is very small. The consequence of these properties introduces interesting features into the quantum heat engine performance, as we will discuss shortly in Sec. III.
@@
where $\beta\!=\!1/(k_{\mathrm{B}}T)$ is the inverse temperature $T$ with the Boltzmann constant $k_{\mathrm{B}}$. It is worth noting that although the upper limit of the sums is formally written as $\infty$, only bound vibrational levels up to $n_{\text{max}}$ contribute; terms beyond vanish identically.In the energy representation, it is straightforward to calculate the partition function using the $q$-deformed Morse potential eigen energies (3), we obtain
@@ III.1 Quantum Carnot Cycle
Similar to the classical Carnot cycle, the cycle is reversible, and the energy eigenvalues must satisfy the reversibility condition for all states [15]. This cycle comprises four strokes, see Fig. 2, namely; isothermal expansion, adiabatic expansion, isothermal compression, and adiabatic compression. The strokes $\text{B}\rightarrow\text{C}$ and $\text{D}\rightarrow\text{A}$ are adiabatic processes, while $\text{A}\rightarrow\text{B}$ and $\text{C}\rightarrow\text{D}$ are isothermal processes. During adiabatic strokes, $\mathcal{P}_{n}(\text{C})\!=\!\mathcal{P}_{n}(\text{B})$, and $\mathcal{P}_{n}(\text{A})\!=\!\mathcal{P}_{n}(\text{D})$; that means $\mathcal{S}(\text{C})\!=\!\mathcal{S}(\text{B})$, and $\mathcal{S}(\text{A})\!=\!\mathcal{S}(\text{D})$. The heat exchanges during the cycle are given as; $\mathcal{Q}_{\text{h}}\!=\!T_{\text{h}}[\mathcal{S}(\text{B})-\mathcal{S}(\text{A})]>0$, and $\mathcal{Q}_{\text{c}}\!=\!T_{\text{c}}[\mathcal{S}(\text{D})-\mathcal{S}(\text{C})]<0$. The total work output reads
@@
with $i\in\{\text{A},\text{B},\text{C},\text{D}\}$, the inverse temperature $\beta\!=\!1/k_{\text{B}}T$, and $k_{\text{B}}$ is the Boltzmann constant. To realize a quantum heat engine, $\mathcal{W}>0$ must be satisfied [60, 61]. Employing the energy eigenvalue, $\mathcal{E}_{n}$, the entropy is evaluated using the approximation $\sum\limits_{n=1}^{\infty}\exp(-\beta\mathcal{E}_{n})\approx\int_{0}^{\infty}\exp(-\beta\mathcal{E}_{n})dn$ and reads
@@
Figure 2: A pictorial representation of a Carnot thermal machine using a $q$-deformed Morse oscillator as a working substance. The engine cycle consists of two adiabatic strokes ($B\to C$ and $D\to A$) and two isothermal strokes, $A\to B$ at $T_{\text{h}}$ and $C\to D$ at $T_{\text{c}}$, with $T_{\text{h}}>T_{\text{c}}$.
Figure 3: Quantum Carnot engine: (a) The heat exchanges of the working substance with the hot reservoir, $\mathcal{Q}_{\text{h}}$ vs $\mathcal{D}_{e}$ and $q$ (b) The heat exchanges of the working substance with the cold reservoir, $\mathcal{Q}_{\text{c}}$ vs $\mathcal{D}_{e}$ and $q$. (c) The Work output $\mathcal{W}$ vs $\mathcal{D}_{e}$ and $q$. Parameters used: $\alpha_{h}=2.236$, $\alpha_{c}=1$, $T_{c}\!=\!2$ and $T_{h}=10$
@@
In Fig. (3), we present the numerical results of the heat exchanges and the total work extracted against the control/deformation parameter ($q/r_{e}$) and the dissociation energy ($D_{e}/r_{e}$). Fig. 3 (a) shows that the input heat is positive, that is $\mathcal{Q}_{\text{h}}>0$ while Fig. 3(b) demonstrate that the output heat is always negative $\mathcal{Q}_{\text{c}}<0$ within the parameters range. In Fig. 3(c), we observe an increase in work output as both $q/r_{e}$ and dissociation energy $D_{e}/r_{e}$ increase. Both parameters enhance the work output but have a stronger effect when they are at their maximum. This demonstrates that the width of the well, controlled by $\alpha$, and the hot bath temperature both play crucial roles in enhancing the work output: the former governs the expansion and compression of the working substance, while the latter determines the thermal population of excited vibrational states during the isothermal process. Thus, considering that the total work output is positive, we conclude that the system operates in the quantum heat engine regime under the conditions studied. The efficiency of the quantum Carnot $q$-deformed Morse oscillator heat engine is 0.8 at temperatures that satisfy the reversibility condition.
@@ III.2 Quantum Otto Cycle
$\mathcal{Q}_{\text{h}}=\int^{B}_{A}\,\mathcal{E}_{n}d\mathcal{P}_{n}=\,\sum_{n}\mathcal{E}_{n}^{h}[\mathcal{P}_{n}(B)-\mathcal{P}_{n}(D)],$ (15)
@@
$\mathcal{Q}_{\text{c}}=-\int^{D}_{C}\mathcal{E}_{n}dP_{n}\,=\sum\limits_{n}\mathcal{E}_{n}^{c}[\mathcal{P}_{n}(D)-\mathcal{P}_{n}(B)],$ (16)
@@ III.2.1 Changing width of the potential well
where the dimensionless parameters $\Lambda_{2}^{\text{hc*}},2\gamma_{1}^{\text{c}},\gamma_{1}^{\text{h}},\text{and}\Lambda_{4}^{ch}$ are defined explicitly in Eq. (35), and (43) of Appendix (B). Using Eq. (17), we obtain the work output as follows:
$\mathcal{W}^{\text{QO}}=\frac{1}{2\sqrt{\pi}}\left(\frac{\Lambda_{10}+\Lambda_{11}\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}}{\sqrt{\beta_{\text{h}}\xi_{\text{h}}^{2}(-p)}\text{$\beta$c}\beta_{\text{h}}\xi_{\text{h}}^{2}\text{erfc}\left(\gamma_{1}^{\text{h}}\right)}-\frac{p\mathcal{R}_{0}}{\mathcal{R}_{1}}\right)$ (21)
@@
Figure 5: Quantum Otto engine with changing width of the potential well: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$, (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ (c) Work output, $\mathcal{W}$, and (d) Efficiency, $\eta$ as a function of $\mathcal{D}_{e}$ and $q$. Parameters used: $\alpha_{h}\!=\!2.236$, $\alpha_{c}\!=\!1$, $T_{c}\!=\!2$ and $T_{h}\!=\!10$.
@@
Figure 6: Quantum Otto cycle with the changing deformation parameter: (a) Heat exchange with the hot reservoir, $\mathcal{Q}_{\text{h}}$ (b) Heat exchange with the cold reservoir, $\mathcal{Q}_{\text{c}}$ (c) Total work output, $\mathcal{W}$ (d) Efficiency $\eta$ as function of $\mathcal{D}_{e}$ and $\alpha$. Parameters used: $q_{c}=0.8$, $q_{h}=1$, $T_{h}=10$, and $T_{c}=2$.
@@ III.2.3 Changing the strength of the dissociation energy
[…]
@@ IV Conclusions
We presented a theoretical proposal for a quantum thermal machine with the $q$-deformed Morse potential as a working medium. We discussed in detail the different regimes of operation for our heat engine. These modes of operation can be modulated by appropriate adjustment of the anharmonic potential parameters. Furthermore, we present analytical expressions for the work done and the efficiency of the heat engine. We find that these figures of merits depend on the variations of the potential parameters. It is also noted that the performance of both the engine can be regulated or controlled by adjusting the deformation parameter and other system parameters. The potential platform for the realization of our model is a laser-cooled trapped ion as a microscopic heat machine, proposed by Gelbwaser-Klimovsky et al. [64]. However, in our context, the trap will be anharmonic, which will be more suitable for trapping a diatomic molecule. The experimental considerations for the realization of a tunable anharmonic trap have been presented by Home et al. [29]. Moreover, the deformation parameter $q$ can be interpreted as an effective control parameter that encodes deviations from the standard Morse potential, which can be engineered. Our findings may inspire several developments in the emerging field of quantum technologies, such as the design of quantum thermal machines based on anharmonic models.